# Electric Vehicle Routing Problem for Emergency Power Supply: Towards Telecom Base Station Relief

### v1

@@ 4.1. Datasets
They include actual data of two different regions222The names of the regions and specific values are omitted for privacy reasons.: the areas of Real-ev-6 and Real-ev-12 are around 46 and 66 km2, respectively. Table 2 organizes the statistics of these datasets. The locations and specifications of nodes and EVs are actual values. Base stations are distributed in the center and charge stations surround it, emulating the actual situation. We assume that the initial batteries of base stations and EVs are 80 % of their capacities.
@@
They are generated based on the distribution of the real datasets. Table 2 organizes their statistics. The locations of nodes are uniformly sampled within the square area km2. The specific values of base stations, charge stations, and EVs are sampled from the distribution of real datasets. The initial batteries of base stations are randomly sampled within 50 - 100 % of the capacity, and those of EVs are set to 80 % of the capacity. Syn-ev-6S is used only for scalability and generalization tests, which is discussed later.
@@ 5. Experimental results
Syn-ev6 (100 samples) | Syn-ev12 (100 samples) | ||||||||||||||||||
= 12h | = 24h | = 12h | = 24h | ||||||||||||||||
Model | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | |||
w/o EVs | - | 20.1 | - | - | - | 33.3 | - | - | - | 20.1 | - | - | - | 33.3 | - | - | |||
Greed | 189 | 17.7 | 37.2 | 1s | 312 | 29.8 | 62.8 | 1s | 192 | 15.5 | 32.9 | 1s | 317 | 26.6 | 56.4 | 2s | |||
Rand (S=12800) | 142 | 15.3 | 32.0 | 1m | 263 | 26.5 | 55.5 | 2m | 157 | 11.8 | 25.1 | 3m | 283 | 21.1 | 45.1 | 6m | |||
Ours (G) | 81 | 12.5 | 25.9 | 1s | 125 | 25.2 | 51.7 | 1s | 78 | 6.63 | 14.0 | 1s | 131 | 18.2 | 37.8 | 2s | |||
Ours (S=1280) | 78 | 12.4 | 25.5 | 51s | 125 | 24.6 | 50.4 | 1m | 75 | 6.41 | 13.6 | 1m | 129 | 17.4 | 36.0 | 3m | |||
Ours (S=12800) | 78 | 12.4 | 25.5 | 9m | 126 | 24.5 | 50.2 | 15m | 74 | 6.38 | 13.5 | 15m | 129 | 17.2 | 35.7 | 28m | |||
Real-ev6 (1 sample) | Real-ev12 (1 sample) | ||||||||||||||||||
= 12h | = 24h | = 12h | = 24h | ||||||||||||||||
Model | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | |||
w/o EVs | - | 8.25 | - | - | - | 17.5 | - | - | - | 11.3 | - | - | - | 25.9 | - | - | |||
Greed | 30 | 5.87 | 19.4 | 1s | 37 | 13.5 | 42.8 | 1s | 88 | 9.34 | 21.7 | 1s | 107 | 20.9 | 47.1 | 1s | |||
Rand (S=12800) | 32 | 3.74 | 13.1 | 1s | 63 | 10.0 | 33.8 | 2s | 120 | 4.98 | 12.8 | 5s | 215 | 14.2 | 34.2 | 8s | |||
Cp-Tsn1.0 | 16 | 5.62 | 17.9 | 1m | 24 | 12.6 | 39.37 | 30m | 47 | 4.54 | 10.62 | 22s | 98 | 14.2 | 32.44 | 30m | |||
Cp-Tsn0.5 | 16 | 3.40 | 11.2 | 11m | 31 | 11.6 | 36.7 | 30m | 53 | 3.32 | 8.06 | 1m | 105 | 15.0 | 34.25 | 30m | |||
Ours (G) | 17 | 2.53 | 8.61 | 1s | 26 | 8.46 | 27.1 | 1s | 59 | 1.06 | 3.26 | 1s | 99 | 11.1 | 25.6 | 1s | |||
Ours (S=1280) | 21 | 2.03 | 7.27 | 1s | 35 | 8.37 | 27.3 | 2s | 61 | 0.65 | 2.41 | 4s | 101 | 10.8 | 25.2 | 6s | |||
@@ 5.3. Generalization Test
The time horizon and the number of EVs/nodes vary depending on the situation. However, it is impracticable to train our solver for each case. Therefore, our solver requires the generalization ability for unseen settings. We here use the same datasets and our solvers as the scalability test. The bottom row of Figure 5 shows the objective values in each dataset, which are evaluated at the same time as the computation time in the scalability test. We report the comparison between our solver and naive approaches (Greed and Rand).555We here exclude Cp-Tsn∗ because it exceeds the index limit of OR-Tools (i.e., outputs IndexError: list index out of range) in the middle of increasing the parameters. Our solver consistently outperforms the naive approaches at every parameter setting, where our solver is trained on the most-left parameter setting in each graph. Notably, we observe that our solver effectively increases the improvement ratio in the unseen settings of 18 and 24 EVs. These results demonstrate that our solver possesses reasonable generalization capabilities for unseen settings. Besides, they confirm that in the previous scalability test, our solver provided not only fast computation but also adequate performance.
@@ 6.2. Electric Vehicle Routing Problem
The Electric VRP (EVRP) is a VRP that additionally considers EV battery consumption by travel and recharging of EV battery. Anticipating the introduction of EVs in logistics, various works related to EVRP have appeared lately (refer to (Kucukoglu et al., 2021; Qin et al., 2021) for more comprehensive reviews on this topic). Here, we focus on the problem formulation. There are many variants of EVRP, where the conventional VRP vehicles are replaced with EVs, and their characteristics are considered. They include the Electric Traveling Salesman Problem with Time Windows (ETSPTW) (Roberti and Wen, 2016), EVRP (Lin et al., 2016), EVRP with Time Windows (EVRPTW) (Schneider et al., 2014; Desaulniers et al., 2016; Keskin and Çatay, 2016), and Electric Pickup and Delivery Problem with Time Windows (EPDPTW) (Goeke, 2019). From more detailed perspective, battery consumption estimation according to the EV’s load has also been studied to make the problem more realistic: linear deterministic (Basso et al., 2019; Xiao et al., 2019; Zhao and Lu, 2019), non-linear deterministic (Goeke and Schneider, 2015; Li et al., 2020), and stochastic estimation (Basso et al., 2021; Pelletier et al., 2019; Soysal et al., 2020; Zhang et al., 2020a). Our formulation, EVRP-EPS, can be considered a variant of the Inventory Routing Problem (IRP) (Lagos et al., 2020), which has not yet been studied in the context of EVRP. That is, the inventory corresponds to the base station battery. Regarding battery consumption estimation in EVRP-EPS, linear models are used, but we found them sufficient based on actual driving data in our setting.
@@ 7. Conclusion and Future work
On the other hand, our solver currently faces two limitations: unbalanced work among EVs and only considering Euclidean distance. Balancing the amount of work (e.g., the travel distance and the count of visiting nodes) among EVs is critical for improving their fairness if the drivers are humans. Moreover, the effectiveness of our method would be reduced if the actual distance between two points differs significantly from the Euclidean distance. In future work, we will address these problems by leveraging route-balancing techniques and a variant of Transformers that considers edge features, which represent the actual travel distance.

### v2

@@ 4.1. Datasets
They include actual data of two different regions222The names of the regions and specific values are omitted for privacy reasons.: the areas of Real-ev-6 and Real-ev-12 are around 46 and 66 km, respectively. Table 2 organizes the statistics of these datasets. The locations and specifications of nodes and EVs are actual values. Base stations are distributed in the center and charge stations surround it, emulating the actual situation. We assume that the initial batteries of base stations and EVs are 80 % of their capacities.
@@
They are generated based on the distribution of the real datasets. Table 2 organizes their statistics. The locations of nodes are uniformly sampled within the square area km. The specific values of base stations, charge stations, and EVs are sampled from the distribution of real datasets. The initial batteries of base stations are randomly sampled within 50 - 100 % of the capacity, and those of EVs are set to 80 % of the capacity. Syn-ev-6S is used only for scalability and generalization tests, which is discussed later.
@@ 5. Experimental results
Syn-ev6 (100 samples) | Syn-ev12 (100 samples) | |||||||||||||||
= 12h | = 24h | = 12h | = 24h | |||||||||||||
Model | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time |
w/o EVs | - | 20.1 | - | - | - | 33.3 | - | - | - | 20.1 | - | - | - | 33.3 | - | - |
Greed | 189 | 17.7 | 37.2 | 1s | 312 | 29.8 | 62.8 | 1s | 192 | 15.5 | 32.9 | 1s | 317 | 26.6 | 56.4 | 2s |
Rand (S=12800) | 142 | 15.3 | 32.0 | 1m | 263 | 26.5 | 55.5 | 2m | 157 | 11.8 | 25.1 | 3m | 283 | 21.1 | 45.1 | 6m |
Ours (G) | 81 | 12.5 | 25.9 | 1s | 125 | 25.2 | 51.7 | 1s | 78 | 6.63 | 14.0 | 1s | 131 | 18.2 | 37.8 | 2s |
Ours (S=1280) | 78 | 12.4 | 25.5 | 51s | 125 | 24.6 | 50.4 | 1m | 75 | 6.41 | 13.6 | 1m | 129 | 17.4 | 36.0 | 3m |
Ours (S=12800) | 78 | 12.4 | 25.5 | 9m | 126 | 24.5 | 50.2 | 15m | 74 | 6.38 | 13.5 | 15m | 129 | 17.2 | 35.7 | 28m |
Real-ev6 (1 sample) | Real-ev12 (1 sample) | |||||||||||||||
= 12h | = 24h | = 12h | = 24h | |||||||||||||
Model | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time | dist | down | obj | time |
w/o EVs | - | 8.25 | - | - | - | 17.5 | - | - | - | 11.3 | - | - | - | 25.9 | - | - |
Greed | 30 | 5.87 | 19.4 | 1s | 37 | 13.5 | 42.8 | 1s | 88 | 9.34 | 21.7 | 1s | 107 | 20.9 | 47.1 | 1s |
Rand (S=12800) | 32 | 3.74 | 13.1 | 1s | 63 | 10.0 | 33.8 | 2s | 120 | 4.98 | 12.8 | 5s | 215 | 14.2 | 34.2 | 8s |
Cp-Tsn | 16 | 5.62 | 17.9 | 1m | 24 | 12.6 | 39.37 | 30m | 47 | 4.54 | 10.62 | 22s | 98 | 14.2 | 32.44 | 30m |
Cp-Tsn | 16 | 3.40 | 11.2 | 11m | 31 | 11.6 | 36.7 | 30m | 53 | 3.32 | 8.06 | 1m | 105 | 15.0 | 34.25 | 30m |
Ours (G) | 17 | 2.53 | 8.61 | 1s | 26 | 8.46 | 27.1 | 1s | 59 | 1.06 | 3.26 | 1s | 99 | 11.1 | 25.6 | 1s |
Ours (S=1280) | 21 | 2.03 | 7.27 | 1s | 35 | 8.37 | 27.3 | 2s | 61 | 0.65 | 2.41 | 4s | 101 | 10.8 | 25.2 | 6s |
@@ 5.3. Generalization Test
The time horizon and the number of EVs/nodes vary depending on the situation. However, it is impracticable to train our solver for each case. Therefore, our solver requires the generalization ability for unseen settings. We here use the same datasets and our solvers as the scalability test. The bottom row of Figure 5 shows the objective values in each dataset, which are evaluated at the same time as the computation time in the scalability test. We report the comparison between our solver and naive approaches (Greed and Rand).555We here exclude Cp-Tsn because it exceeds the index limit of OR-Tools (i.e., outputs IndexError: list index out of range) in the middle of increasing the parameters. Our solver consistently outperforms the naive approaches at every parameter setting, where our solver is trained on the most-left parameter setting in each graph. Notably, we observe that our solver effectively increases the improvement ratio in the unseen settings of 18 and 24 EVs. These results demonstrate that our solver possesses reasonable generalization capabilities for unseen settings. Besides, they confirm that in the previous scalability test, our solver provided not only fast computation but also adequate performance.
@@ 6.2. Electric Vehicle Routing Problem
The Electric VRP (EVRP) is a VRP that additionally considers EV battery consumption by travel and recharging of EV battery. Anticipating the introduction of EVs in logistics, various works related to EVRP have appeared lately (refer to (Kucukoglu et al., 2021; Qin et al., 2021) for more comprehensive reviews on this topic). Here, we focus on the problem formulation. There are many variants of EVRP, where the conventional VRP vehicles are replaced with EVs, and their characteristics are considered. They include the Electric Traveling Salesman Problem with Time Windows (ETSPTW) (Roberti and Wen, 2016), EVRP (Lin et al., 2016), EVRP with Time Windows (EVRPTW) (Schneider et al., 2014; Desaulniers et al., 2016; Keskin and Çatay, 2016), and Electric Pickup and Delivery Problem with Time Windows (EPDPTW) (Goeke, 2019). From more detailed perspective, battery consumption estimation according to the EV’s load has also been studied to make the problem more realistic: linear deterministic (Basso et al., 2019; Xiao et al., 2019; Zhao and Lu, 2019), non-linear deterministic (Goeke and Schneider, 2015; Li et al., 2020), and stochastic estimation (Basso et al., 2021; Pelletier et al., 2019; Soysal et al., 2020; Zhang et al., 2020a). Our formulation, EVRP-EPS, can be considered a variant of the Continuous Inventory Routing Problem (CIRP) (Lagos et al., 2020). However, it has not yet been studied in the context of EVs for emergency power supply, where the inventory corresponds to the base station battery. Regarding battery consumption estimation in EVRP-EPS, linear models are used, but we found them sufficient based on actual driving data in our setting.
@@ 7. Conclusion and Future work
On the other hand, our solver currently faces two limitations: unbalanced work among EVs and only considering Euclidean distance. Balancing the amount of work (e.g., the travel distance and the count of visiting nodes) among EVs is critical for improving their fairness if the drivers are humans. Moreover, the effectiveness of our method would be reduced if the actual distance between two points differs significantly from the Euclidean distance. In future work, we will address these problems by leveraging route-balancing techniques and a variant of Transformers that considers edge features, representing the actual travel distance.
Acknowledgement
In the illustrations, we used free materials from Flaticon.com (https://www.flaticon.com/).